Suppose that inflation rates in the U.S. and France are expected to be, on average, annually 7.5% and 7%, respectively, over the next five years. If the current spot rate is $3/FRF, then what is the expected spot value of the franc (USD/FRF) in five years?

Select one:

a. USD3.014/FRF

b. USD2.931/FRF

c. USD3.028/FRF

d. USD3.071/FRF

e. USD0.341/FRF
Answer #1

As inflation in US is higher, USD depreciate or FRF will appreciate,
Expected Spot=Current Spot*((1+inflation in US)/(1+inflation in France))^t=3*(1.075/1.07)^5=3.071